# On the Utility of Accounting for Human Beliefs about AI Behavior in Human-AI Collaboration

### v1

@@
Abstract
To enable effective human-AI collaboration, merely optimizing AI performance while ignoring humans is not sufficient. Recent research has demonstrated that designing AI agents to account for human behavior leads to improved performance in human-AI collaboration. However, a limitation of most existing approaches is their assumption that human behavior is static, irrespective of AI behavior. In reality, humans may adjust their action plans based on their observations of AI behavior. In this paper, we address this limitation by enabling a collaborative AI agent to consider the beliefs of its human partner, i.e., what the human partner thinks the AI agent is doing, and design its action plan to facilitate easier collaboration with its human partner. Specifically, we developed a model of human beliefs that accounts for how humans reason about the behavior of their AI partners. Based on this belief model, we then developed an AI agent that considers both human behavior and human beliefs in devising its strategy for working with humans. Through extensive real-world human-subject experiments, we demonstrated that our belief model more accurately predicts humans’ beliefs about AI behavior. Moreover, we showed that our design of AI agents that accounts for human beliefs enhances performance in human-AI collaboration.
1 Introduction
@@
In particular, optimizing the AI system in isolation is not sufficient to enhance the performance of human-AI collaboration. Both the human and the AI agent need to coordinate by inferring the goals and intentions of their counterpart and taking complementary actions. For example, in AI-assisted decision-making, researchers have demonstrated that instead of optimizing AI by itself, training AI to focus on improving in areas where humans typically struggle can significantly enhance the performance of human-AI teams (Bansal et al., 2021; Wilder et al., 2021). In human-AI collaboration, Carroll et al. (2019) show that incorporating models of human behavior into the training of the AI agent leads to higher collaborative performance compared to training the agent to play with themselves through self-play (Silver et al., 2018). These studies highlight the importance of integrating human behavior into the design of collaborative AI. However, a key limitation of these research efforts is that they mostly assume that human behavior remains static, irrespective of the actions and behavior of the AI counterpart. In practice, humans may modify their behavior in response to their beliefs about what AI agents intend to do based on their observations of AI behavior.
Figure 1: Example task.
In this work, we argue that, in addition to incorporating human behavior, designing an AI agent that account for humans’ beliefs about AI behavior could significantly improve their collaborative performance. Consider a simple illustrative example in Figure 1, where the human and AI agents need to go to both goals $\{G1,G2\}$ to complete the task. Assume that the human prefers $G2$ over $G1$ because it requires fewer steps. When designing a collaborative AI that accounts for human behavior only, the AI would choose to go to $G1$, leading to lower overall collaborative performance. However, if the human would adjust their behavior based on their belief about AI behavior (e.g., they will avoid going to the same goal that the AI is going towards), we could incorporate this knowledge into the design of the AI. For example, the AI can choose to go to $G2$, anticipating that the human will account for that and go to $G1$, leading to improved overall performance. This example underscores the potential benefits of incorporating human beliefs about AI behavior in AI design. Meanwhile, it also highlights the importance of developing accurate models of human beliefs and incorporating them into the AI design.
@@
In this work, we aim to examine the benefits of incorporating human beliefs about AI behavior in the design of AI to improve human-AI collaboration. Our work focuses on the setting where the human and AI agents need to coordinate their goals (i.e., they need to complete different subtasks in order to accomplish the overall task). More formally, we formulate the human-AI collaboration environment as a multi-player goal-oriented Markov decision process (MDP). We begin our investigation by developing models of human beliefs regarding AI behavior. This modeling effort extends the level-$k$ reasoning framework (Stahl II and Wilson, 1994) to account for suboptimal human behavior. Specifically, we first develop a behavioral level-0 model that assumes agents take actions without considering the behavior of other agents. We then enhance the belief model by introducing a behavioral level-1 model, which assumes humans interpret the behavior of another player as if the other agent adheres to the behavioral level-0 model. With the models of human behavior and beliefs established, we proceed to develop collaborative AI agents that incorporate different assumptions about human models.
@@
To examine our proposed approaches, we conduct extensive human-subject experiments on two environments. To develop and assess the model of human behavior (i.e., the behavioral level-00 model), we first train models of human behavior using behavioral cloning on real-world human behavior and evaluate their performance. For the development and assessment of the model of human beliefs (i.e., the behavioral level-$1$ model), we conduct experiments that present each participant with a trace of behavior by an agent and ask them to infer the agent’s goal, examining whether our model leads to accurate predictions of human inference on the agent’s goal. Moreover, based on the belief model, we also explore the possibility of developing AI policies such that humans can more easily infer the goals of AI from its actions. Finally, utilizing the developed models of human behavior and beliefs, we conduct a final set of human-subject experiments that pair each human participant with an AI agent in a two-player coordination game. We assess the human-AI collaborative performance for different designs of AI agents. Overall, our results demonstrate the effectiveness of our developed model of human behavior and beliefs. Furthermore, we show that AI agents accounting for models of human behavior and beliefs achieve better collaborative performance with humans, compared to AI agents that do not consider human beliefs or those that disregard both human behavior and beliefs.
Contributions. The main contributions of our work can be summarized as follows:
@@
    We develop models of human beliefs, inspired by the level-$k$ framework, that account for human suboptimality. Moreover, we have designed AI agents that can take explicable action plans, making it easier for humans to infer AI goals from observing AI actions.
@@
    We have designed collaborative AI agents, considering both models of human behavior and human beliefs about AI behavior, aiming to optimize the performance of human-AI collaboration.
@@
    Through extensive human-subject experiments with 1,690 participants, we demonstrate that our belief models are effective and can help us design AI policies that are explicable to humans. Moreover, our design of collaborative AI that incorporates human beliefs about AI behavior improves the collaborative performance when paired with real-world human participants. The results demonstrate both the importance of incorporating human beliefs about AI behavior in the design of collaborative AI and the effectiveness of our approach.
2 Related Work
@@
Our work joins the recent growing research in human-AI collaborations (Carroll et al., 2019; Jaderberg et al., 2019; Bansal et al., 2021; Wilder et al., 2021). Bansal et al. (2021) demonstrate that optimizing AI in isolation may lead to suboptimal performance for human-AI collaboration. Wilder et al. (2021) show that training AI systems to complement humans, performing better in areas where humans struggle, leads to improved collaborative performance. Carroll et al. (2019) illustrate that incorporating a human model, learned from human data, into the training of AI results in enhanced performance when these AI systems work with real humans. Additionally, Jaderberg et al. (2019) use population-based reinforcement learning to improve the robustness of trained AI agents. Our work extends this line of work by proposing to incorporate not only human behavior but also human beliefs of AI behavior into the design of collaborative AI.
@@
From a technical perspective, our work involves understanding and modeling humans in decision-making. There has been a significant amount of work in the literature on modeling human behavior. For example, Inverse Reinforcement Learning (IRL)(Ng et al., 2000; Abbeel and Ng, 2004; Ramachandran and Amir, 2007) aims to infer the reward functions in Markov Decision Processes (MDPs) through observing demonstrations of the optimal policy. If the demonstrator is a human being, the demonstrations could be noisy or contain behavioral biases. Studies (Evans et al., 2016; Shah et al., 2019; Hughes et al., 2020; Zhi-Xuan et al., 2020) have aimed to incorporate human behavioral biases in the inference process and infer both the rewards and biases simultaneously. Imitation Learning (Hussein et al., 2017) also aims at developing models that can mimic human behavior from demonstrations. Recently, there have also been an increasing amount of research efforts that incorporates human models in computational and machine learning frameworks Kleinberg and Oren (2014); Tang and Ho (2019, 2021); Masters et al. (2021a, b); Yu and Ho (2022); Yu et al. (2023); Feng et al. (2024).
@@
From the perspective of modeling human beliefs over others’ behavior, this has been discussed in the literature of level-$k$ reasoning (Stahl II and Wilson, 1994; Gill and Prowse, 2016) in economics and theory of mind (Premack and Woodruff, 1978; Westby and Riedl, 2023). As a few examples, Yu et al. (2006) have found that real human behavior is close to level-$1$ and level-$2$ reasoning models in cooperative games. Albrecht and Stone (2018) survey works on autonomous agents modeling the beliefs and intentions of other agents. They distinguish methods for modeling stationary or changing agent behaviors. The belief model includes theory of mind, recursive reasoning models, plan recognition, partially observable Markov decision processes (POMDPs), and others. Baker et al. (2011) propose a Bayesian theory of mind to describe human modeling of joint belief of state in a partially observable MDP and conduct human experiments in (Baker et al., 2009, 2017). A similar human modeling approach is adopted by Wu et al. (2021) and applied to the Overcooked experiment. Their idea is to split delivering a dish into several subtasks, such as picking up an ingredient or putting an ingredient into a pot, and the player is inferring others’ subtasks and selecting their own corresponding subtask before taking actions.
3 Problem Formulation and Methods
@@
3.1 Decision-Making Environment
@@
We formulate the human-AI cooperative decision-making environment as a multi-agent Markov decision process (MDP), represented by $W=\langle S,\alpha,{A_{i}|_{i\in[\alpha]}},R,P\rangle$. $S$ is a finite set of states; $\alpha$ is a finite set of players; $A_{i}$ is the action set available to player $i$. $P:S\times A_{1}\times\dots\times A_{|\alpha|}\times S\rightarrow[0,1]$ is the transition function that determines the next state given all players’ joint actions. $R:S\times A_{1}\times\dots\times A_{|\alpha|}\rightarrow\mathbb{R}$ is the joint reward function assigned to all players given their joint actions. Note that in this work we consider the cooperative setting. Therefore, our formulation only incorporates a single reward function $R$ for all players, though this can be easily extended.
@@
While our formulation could address cases with multiple human and AI players, in this work, we focus on a two-player cooperative game, i.e., $\alpha=\{1,2\}$, where one player is a human and the other is the AI. During the decision-making process, neither the human nor the AI knows the other player’s next action or future plans, and they cannot communicate directly. However, they can observe each other’s past actions, enabling them to infer about the other player and modify their own actions accordingly.
@@
Goal-oriented MDP. In this work, we focus on the setting with goal-oriented MDP, where there is a set of goals $G=\{g_{1},\dots,g_{k}\}\subseteq S$, which are subset of states that are terminal states, i.e., $P(g|g,a)=1~{}\forall g\in G,a\in A$. Moreover, the decision-making agent only receives rewards when arriving at one of the goal states.
3.2 Modeling Human Behavior and Human Beliefs about AI Behavior
@@
Our models are motivated by the level-$k$ framework (Stahl II and Wilson, 1994) in economics. In particular, we start by considering humans as level-00 agents that do not account for others’ behavior in the environment. Differing from the literature, we address the natural situation that level-00 agents do not behave optimally, and we call this model behavioral level-00 agents. To account for human beliefs, [^1] In this work, human beliefs refer to humans’ belief about the goal of the AI agent. we model humans as an extension of the level-$1$ agents, that assume other agents are behavioral level-00 agents [^2] Our model can iteratively progress to higher level-$k$ agents. However, we focus on the case with $k\leq 1$. and update their beliefs in a Bayesian manner based on the observations of others’ behavior.
@@
Modeling human behavior. We first model human behavior under the assumption that humans do not consider other players in the environment (or that they consider other players as a part of the environment without strategically responding). In this case, a human behavior model can be represented as $H:W\rightarrow\Pi$, mapping a given environment $w\in W$ to a policy $\pi=H(w)$. We give two examples of human behavior models utilized in our work below.
@@
    Behavioral level-00 model. We also consider the case that we can learn human behavioral models from their historical behavioral traces through behavioral cloning (Pomerleau, 1988; Torabi et al., 2018). Behavioral cloning is one of the imitation learning approaches, which learns a policy from human demonstration by building a map from states to actions with supervised learning methods Bain and Sammut (1995). We build a fully connected neural network, where the input is the state encoding, and output is the probability over action space, and train the model with standard gradient descent method with cross-entropy loss.
Modeling human beliefs. We now describe how we develop models for human beliefs. As summarized earlier, we consider cases where human decision-makers assume that other agents in the environments are behavioral level-00 agents. In our discussion, we also describe this belief model as behavioral level-$1$ agents.
[…]
$B(g|(s,a)_{1:t})\propto\lambda(g)Pr((s,a)_{1:t}|g)=\lambda(g)\prod_{1:t}\pi(a_{i}|s_{i},g)P(s_{i}|s_{i-1},a_{i-1})$ (1)
3.3 Designing Collaborative AI Agents
To illustrate the effectiveness of incorporating models of human beliefs, we train different AI agents that work with humans in human-AI collaboration via simulations and real-world human subject experiments.

### v2

# On the Utility of Accounting for Human Beliefs about AI Intention in Human-AI Collaboration

@@
Abstract.
To enable effective human-AI collaboration, merely optimizing AI performance without considering human factors is insufficient. Recent research has shown that designing AI agents that take human behavior into account leads to improved performance in human-AI collaboration. However, a limitation of most existing approaches is their assumption that human behavior remains static, regardless of the AI agent’s actions. In reality, humans may adjust their actions based on their beliefs about the AI’s intentions, specifically, the subtasks they perceive the AI to be attempting to complete based on its behavior. In this paper, we address this limitation by enabling a collaborative AI agent to consider its human partner’s beliefs about its intentions, i.e., what the human partner thinks the AI agent is trying to accomplish, and to design its action plan accordingly to facilitate more effective human-AI collaboration. Specifically, we developed a model of human beliefs that captures how humans interpret and reason about their AI partner’s intentions. Using this belief model, we created an AI agent that incorporates both human behavior and human beliefs when devising its strategy for interacting with humans. Through extensive real-world human-subject experiments, we demonstrate that our belief model more accurately captures human perceptions of AI intentions. Furthermore, we show that our AI agent, designed to account for human beliefs over its intentions, significantly enhances performance in human-AI collaboration.
Key words and phrases:
Human-AI Collaboration, Reinforcement Learning, Belief Inference
footnotetext: Preprint. Under review.
1. Introduction
@@
In particular, optimizing the AI system in isolation is often shown to be insufficient for enhancing the performance of human-AI collaboration. To optimize collaborative performance, AI agents need to account for their human counterparts when deciding on their strategies. For example, in AI-assisted decision-making, researchers have demonstrated that, instead of optimizing AI independently, training AI to focus on improving in areas where humans typically struggle can significantly enhance the performance of human-AI teams (Bansal et al., 2021; Wilder et al., 2021). In human-AI collaboration, Carroll et al. (2019) show that incorporating models of human behavior into the training of AI agents leads to higher collaborative performance compared to training the agent through self-play (Silver et al., 2018). These studies highlight the importance of integrating human behavior into the design of collaborative AI. However, a key limitation of these research efforts is that they mostly assume human behavior remains static, irrespective of the actions and behavior of the AI counterpart. In practice, humans may modify their behavior in response to their beliefs about the intentions of AI agents based on their observations of AI behavior.
In this work, we argue and aim to demonstrate that, in addition to incorporating human behavior, designing an AI agent that accounts for humans’ beliefs about AI intentions can significantly improve their collaborative performance. To illustrate this argument, consider a simple example task in Figure 1. In this task, both the human and AI need to collaborate to put out a fire in the center of the map. To successfully extinguish the fire, they need to use both a fire extinguisher and a hydrant, with the human and AI agent each retrieving one of them. Assuming they cannot directly communicate but can observe the direction the other is moving towards, if the AI agent believes the human will simply take the tool closest to them (the hydrant) and will not adjust their actions based on the AI’s behavior, the AI would choose to take the longer route to get the fire extinguisher. However, if the human would actually update their actions based on their inference of what the AI is trying to do from observing the AI’s actions, the more efficient strategy for the AI would be to move towards the hydrant and let the human grab the fire extinguisher. This example task demonstrates that, if humans would update their behavior based on their beliefs over AI intentions, we can design AI that account for human belief to improve overall human-AI performance.
(a) When the AI agent does not account for human beliefs about AI intentions and only considers human behavior, it would assume that the human would grab the closer hydrant. Therefore, the AI would go for the fire extinguisher, which would take 12 steps to put out the fire.
(b) When the AI agent accounts for human beliefs over AI intentions, it will try to reveal its intentions for humans so humans can take a complementary tool that would reach the overall better outcome. In this case, it takes 9 steps for the fire to be put out.
Figure 1. An example human-AI cooperation task that demonstrates the benefits for the AI to account for human beliefs about AI intentions.
In this work, we investigate and examine the benefits of incorporating human beliefs about AI intentions in the design of AI to improve human-AI collaboration. To operationalize the concept of AI intentions, our work focuses on a setting where human and AI agents need to coordinate their goals (e.g., they need to complete different subtasks) to accomplish the overall task. In this setting, the intention of the AI is defined as the goal or subtask the AI is attempting to achieve.
@@
More formally, we formulate the human-AI collaboration environment as a multi-player goal-oriented Markov decision process (MDP). We begin our investigation by developing models of human beliefs regarding AI intentions. This modeling effort extends the level-$k$ reasoning framework (Stahl II and Wilson, 1994) to account for suboptimal human behavior. Specifically, we first develop a behavioral level-0 model that assumes agents take actions without considering the behavior of other agents. We then develop the belief model by introducing a behavioral level-1 model, which assumes humans interpret the behavior of another player as if the other agent take actions according to the behavioral level-0 model. With the models of human behavior and beliefs established, we proceed to develop collaborative AI agents that incorporate different assumptions about human models.
@@
To examine our proposed approaches, we conduct extensive human-subject experiments. To develop and assess the model of human behavior (i.e., the behavioral level-00 model), we first train models of human behavior using behavioral cloning on real-world human data and evaluate their performance. For the development and assessment of the model of human beliefs about AI intentions (i.e., the behavioral level-$1$ model), we conduct experiments that present each participant with a trace of behavior by an agent and ask them to infer the agent’s goal, examining whether our model accurately predicts human inference of the agent’s intention. Moreover, based on the belief model, we explore the possibility of developing AI policies that make it easier for humans to infer the intentions of AI from its actions. Finally, utilizing the developed models of human behavior and beliefs about AI intentions, we conduct a final set of human-subject experiments in which each human participant pairs with an AI agent in a two-player coordination game. We assess the human-AI collaborative performance for different designs of AI agents. Overall, our results demonstrate the effectiveness of our developed models of human behavior and beliefs about AI intentions. Furthermore, we show that AI agents accounting for models of human behavior and beliefs about AI intentions achieve better collaborative performance when working with humans compared to AI agents that either do not consider human beliefs or disregard both human behavior and beliefs.
Contributions. In this work, we demonstrate that incorporating human beliefs about AI intentions in the design of collaborative AI can significantly improve human-AI collaborative performance. More concretely, the main contributions of our work are summarized as follows:
@@
    We develop models of human beliefs about AI intentions, inspired by the level-$k$ framework, and extend this framework to account for human suboptimality. In addition, we design AI agents capable of executing explicable action plans, making it easier for humans to infer AI intentions based on observed AI actions.
@@
    We design collaborative AI agents that incorporate both models of human behavior and human beliefs about AI intentions, aiming to optimize the performance of human-AI collaboration.
@@
    Through extensive human-subject experiments, we demonstrate that our belief models are effective and can guide the design of AI policies that are explicable to humans. Moreover, our design of collaborative AI that integrates human beliefs about AI intentions significantly improves collaborative performance when paired with real-world human participants. These results underscore both the importance of incorporating human beliefs about AI intentions in the design of collaborative AI and the effectiveness of our approach.
2. Related Work
@@
Our work contributes to the growing body of research on human-AI collaboration (Carroll et al., 2019; Jaderberg et al., 2019; Bansal et al., 2021; Wilder et al., 2021). For example, Bansal et al. (2021) demonstrate that optimizing AI in isolation may lead to suboptimal performance in human-AI collaboration, compared to training AI with consideration for human performance. Wilder et al. (2021) show that training AI systems to complement humans by performing better in areas where humans struggle results in improved collaborative performance. Carroll et al. (2019) illustrate that incorporating a human model, learned from human data, into the training of AI enhances performance when these AI systems collaborate with real humans. Our work extends this line of research by not only accounting for human behavior when designing AI but also incorporating human beliefs about AI intentions.
While these approaches require knowledge or data of humans, there have been zero-shot coordination techniques being proposed to train cooperative AI agents when human data are not available. Jaderberg et al. (2019) use population-based reinforcement learning to improve the robustness of trained AI agents. Strouse et al. (2021) develop a variety of self-play agents by varying random seeds of neural network initialization and various checkpoints, and train a best response model tailored to the agent population, discovering that humans prefer to partner with this model. Zhao et al. (2023) utilize the advances of maximum entropy reinforcement learning to encourage the diversity of agent populations and show the learned AI models could adapt to real humans. In this work, we have included these approaches as baselines and show that accounting for human beliefs lead to better performance in our setting.
@@
From a technical perspective, our work focuses on understanding and modeling human decision-making. A significant body of literature has explored modeling human behavior. For example, inverse reinforcement learning (IRL)(Ng et al., 2000; Abbeel and Ng, 2004; Ramachandran and Amir, 2007) aims to infer reward functions in Markov decision processes (MDPs) by observing demonstrations of the optimal policy. When the demonstrator is human, these demonstrations may be noisy or exhibit behavioral biases. Several studies(Evans et al., 2016; Shah et al., 2019; Hughes et al., 2020; Zhi-Xuan et al., 2020) have addressed this by incorporating human behavioral biases into the inference process, aiming to infer both the rewards and biases simultaneously. Imitation Learning (Hussein et al., 2017) also seeks to develop models that can mimic human behavior based on demonstrations. Recently, there has been a growing amount of research that incorporates human models into computational and machine learning frameworks (Kleinberg and Oren, 2014; Tang and Ho, 2019, 2021; Masters et al., 2021a, b; Yu and Ho, 2022; Yu et al., 2023; Feng et al., 2024).
@@
From the perspective of modeling human beliefs about others’ intentions, this has been discussed in the literature on level-$k$ reasoning (Stahl II and Wilson, 1994; Gill and Prowse, 2016) in economics and in the theory of mind (Premack and Woodruff, 1978; Westby and Riedl, 2023). For instance, Yu et al. (2006) found that real human behavior aligns closely with level-$1$ and level-$2$ reasoning models in cooperative games. Albrecht and Stone (2018) survey works on autonomous agents modeling the beliefs and intentions of other agents, distinguishing between methods for modeling stationary or changing agent behaviors. These belief models include theory of mind, recursive reasoning models, plan recognition, partially observable Markov decision processes (POMDPs), and others. Baker et al. (2011) propose a Bayesian theory of mind to describe how humans model joint beliefs of state in a partially observable MDP, conducting human experiments in the works of Baker et al. (2009, 2017). Similarly, Wu et al. (2021) apply a human modeling approach in the Overcooked experiment, breaking down the task of delivering a dish into several subtasks, such as picking up an ingredient or placing an ingredient into a pot, where the player infers others’ intentions regarding these subtasks and selects their own accordingly before taking actions. Our belief inference framework differs from previous theory of mind literature by modeling and incorporating human beliefs about AI intentions specifically, using a more realistic human model as the basis for inferring these intentions.
3. Problem Formulation and Methods
@@
3.1. Decision-Making Environment
@@
We formulate the human-AI cooperative decision-making environment as a multi-agent MDP, represented by $W=\langle S,\alpha,{A_{i}|_{i\in[\alpha]}},R,P\rangle$, where $S$ is a finite set of states, $\alpha$ is a finite set of players, $A_{i}$ is the action set available to player $i$, $P:S\times A_{1}\times\dots\times A_{|\alpha|}\times S\rightarrow[0,1]$ is the transition function that determines the next state given all players’ joint actions, and $R:S\times A_{1}\times\dots\times A_{|\alpha|}\rightarrow\mathbb{R}$ is the joint reward function assigned to all players given their joint actions. Note that in this work we consider the cooperative setting. Therefore, our formulation only incorporates a single reward function $R$ for all players, though this can be easily extended.
@@
While our formulation could address cases with multiple human and AI players, in this work, we focus on a two-player cooperative game, i.e., $\alpha={1,2}$, where one player is a human and the other is the AI. During the decision-making process, neither the human nor the AI knows the other player’s next action or future plans, and they cannot communicate directly. However, they can observe each other’s past actions, allowing them to infer the intentions of the other player and adjust their own actions accordingly.
@@
Agent intentions. In this work, we focus on a setting with a goal-oriented MDP, where there is a set of goals $G={g_{1},\dots,g_{k}}\subseteq S$. The decision-making agent only receives rewards when arriving at one of the goal states, and the goal states are terminal states, i.e., $P(g|g,a)=1\leavevmode\nobreak\ \forall g\in G,a\in A$. Both humans and AI agents can observe past actions but not the specific goals the other agents are trying to achieve. We utilize this formulation to operationalize agent intentions; that is, the goal an agent is trying to reach is defined as the agent’s intention. In the discussion of this paper, we interchangeably use both goals and intentions to refer to intentions.
3.2. Modeling Human Beliefs about AI Intention
@@
Our models are motivated by the level-$k$ framework (Stahl II and Wilson, 1994) in economics. In particular, we start by considering humans as level-00 agents who do not account for others’ behavior in the environment. Unlike previous literature, we address the realistic scenario where level-00 agents may not behave optimally, and we call this model behavioral level-00 agents. To account for human beliefs about AI intentions, [^1] In this work, we sometimes use human beliefs to refer specifically to humans’ beliefs about the intention of the AI agent. we model humans as an extension of level-$1$ agents, who assume other agents are behavioral level-00 agents [^2] Our model can iteratively extend to higher level-$k$ agents; however, we focus on the case with $k\leq 1$. and update their beliefs about AI intentions in a Bayesian manner based on the observation of others’ behavior.
@@
Modeling human behavior. We first describe our models of human behavior under the assumption that humans do not consider other players in the environment (or that they view other players as part of the environment without strategically responding). Our approach to modeling human behavior follows standard methods in the literature and primarily serves as a foundation for developing models of human beliefs. In particular, a human behavior model can be represented as $H:W\rightarrow\Pi$, mapping a given environment $w\in W$ to a policy $\pi=H(w)$. We provide two examples of human behavior models used in our work below.
@@
    Behavioral level-00 model. We also consider the case where we can learn human behavioral models from their historical behavioral traces using behavioral cloning approaches (Pomerleau, 1988; Torabi et al., 2018). Behavioral cloning is an imitation learning approach that learns a policy from human demonstrations by mapping states to actions using supervised learning methods (Bain and Sammut, 1995). We build a fully connected neural network, where the input is the state encoding and the output is the probability distribution over the action space. The model is trained using the standard gradient descent method with a cross-entropy loss.
Modeling human beliefs. We now describe how we develop models for human beliefs about AI intentions. As summarized earlier, our belief models follows the ideas of level-$k$ reasoning and assume that human decision-makers perceive other agents in the environment as behavioral level-00 agents. In our discussion, we also refer to this belief model as behavioral level-$1$ agents.
More specifically, we model the human belief updating process as a Bayesian inference process, which can be represented as in Equation (1), where $\lambda(g)$ represents the human’s prior belief distributions of AI intentions, and $Pr(s_{t},a_{t}|g)$ denotes the probability of observing $(s_{t},a_{t})$ given the intention (goal) $g$, according to the policy model $\pi(a|s,g)$. This policy model represents how humans perceive the actions of other players. If a human believes the other agent is following the standard model, then the policy is derived from the optimal policy. On the other hand, if humans believe the agent follows the behavioral cloning model, then the policy will be the output of the human model learned from data.
$B(g|(s,a)_{1:t})$ $\propto\lambda(g)Pr((s,a)_{1:t}|g)$ (1)
$=\lambda(g)\prod_{1:t}Pr(s_{i},a_{i}|g)$
$=\lambda(g)\prod_{1:t}\pi(a_{i}|s_{i},g)P(s_{i}|s_{i-1},a_{i-1})$
3.3. Designing Explicable AI Policy
Prior research (Ullman et al., 2009; Baker et al., 2011; Zhi-Xuan et al., 2020) has demonstrated that humans often find it challenging to infer the intentions of other agents based on their behavioral traces, which also aligns with our observations in human-subject experiments (see Section 4 and the appendix for comprehensive results). Therefore, before we delve into the design of collaborative AI, we also explore whether we can leverage the model of human beliefs about AI intentions to design an explicable AI policy, where the AI takes actions that make it easier for humans to infer its intentions.
To train an explicable AI policy, we incorporate the models of human beliefs into the AI training process. Instead of solely aiming to maximize rewards, we also focus on aligning the AI policy to increase the probability that humans can infer AI intentions based on AI actions, leveraging the human belief model to account for the human inference process. More specifically, we utilize self-play to train models. In addition to the original MDP rewards, our approach includes additional bonus rewards when the goal inference based on the human belief model is consistent with the specified goal. Specifically, we set the bonus reward to be proportional to the log likelihood of the belief model inference, $\log(\pi(a|s,g))$. With this setting, the trained AI policy would aim to reach the goal while also maximizing the likelihood of revealing its goals, where the likelihood of the goal inference follows the belief model.
3.4. Designing Collaborative AI Agents
We now describe how we design collaborative AI agents that account for different models of human behavior and beliefs. Later, in the experiments, we illustrate the effectiveness of incorporating models of human beliefs by examining the performance of AI agents paired with humans in human-AI collaboration through simulations and real-world human-subject experiments.